OpenStudy (anonymous):

In this next example, why do we add "A" to the RHS and not the LHS. I'm trying to understand nonhomogenous DE...sigh :-( \[y''+4y=e^{3x}\] ... \[y=Ae^{3x}\] I feel like I'm just blindly following a pattern, but I wouldn't be able to recreate this on my own unless I memorize this pattern.

OpenStudy (anonymous):

you're not adding it. it's an arbitrary constant.

OpenStudy (anonymous):

I guess I'm having a hard time expressing myself. I mean why is the A and such on the RHS and not the LHS?...What would I think of when I think of a nonhomogenous DE...what's my goal? Why am I deriving the RHS 2 times and then solving for A?

OpenStudy (anonymous):

You're trying to find what function, y, makes the differential equation true.

OpenStudy (anonymous):

But it seems like something's missing here . . .

OpenStudy (anonymous):

would that mean that \[4y=Ae^{3x}\] \[y'=3Ae^{3x}=0\] \[y''=9Ae^{3x}=1\] ?

OpenStudy (anonymous):

Yes, something like that. If y=Ae^{3x}, then y' = 3Ae^{3x}, and y'' = 9Ae^{3x} Putting that into the original equation, 9Ae^{3x} +4(3Ae^{3x}) = e^{3x}

OpenStudy (anonymous):

Yep. Do we then solve for A?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

this will allow us to get the equation y=.... which was our goal?

OpenStudy (anonymous):

Yes, that will give what is called a particular solution. You'll need initial values / boundary conditions to get an exact solution for those conditions.

OpenStudy (anonymous):

Do you know how to use the characteristic equation to find the complementary solution?

OpenStudy (anonymous):

I guess it's coming together now. but \[y=e^{\alpha x}(c_1cos\beta x+c_2 sin\beta x)\] because the auxiliary equation is r==/- 2i

OpenStudy (anonymous):

I meant +/-2i

OpenStudy (anonymous):

auxiliary is \[r^2+4=0\]

OpenStudy (anonymous):

initial values / boundary conditions....what r_1 and r_2 be considered?

OpenStudy (anonymous):

Right. From the LHS you can treat it as a homogeneous equation to find the characteristic/auxilery, then combine that with the particular solution to get the general solution. Only then do you apply the initial conditions to determine your constants.

OpenStudy (anonymous):

I'm sorry but what are my initial conditions?

OpenStudy (anonymous):

did I ask a really bad question? :S

OpenStudy (anonymous):

I don't know what they are. Do you have any? If you don't have any, then you leave it in general form and you're done (you'll still have undetermined coefficients, but if you don't know initial conditions, you just have to leave it that way.

OpenStudy (anonymous):

Yeah I don't think there were any initial conditions. Ok I think I kinda get nonhomogenous DE's a little better now. I'll try a problem and we'll see how I do. I'll post it in a second. Thanks soo much for all of your help and patience :-)

OpenStudy (anonymous):

I can show you an example of an initial value problem if you like.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Ok, I'll do a simple one so it'll be easy to follow. Given: \[y''-4y'-12y=3e^{5t}, \space y(0)=\frac{18}{7}, \space y'(0)=-\frac{1}{7}\]

OpenStudy (anonymous):

The general solution is: \[y(t)=y_c(t)+Y_p(t)\] Where y_c is the characteristic equation, and Y_p is the particular.

OpenStudy (anonymous):

Sorry *complementary, not "characteristic"

OpenStudy (anonymous):

The characteristic equation is found from the LHS: \[r^2-4r-12=0\] The solutions for r are {-2, 6}

OpenStudy (anonymous):

This makes the complementary equation \[y_c(t)=c_1e^{-2t}+c_2e^{6t}\]

OpenStudy (anonymous):

(This should all look familiar, I hope)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Ok. The particular solution is a guess that y is something like e^x in order to be differentiated and still end up with something like e^x on the RHS, so we guess: \[Y_p(t)=Ae^{5t}\]

OpenStudy (anonymous):

Differentiate that once, then again, and plug y, y', and y'' into the original equation to solve for A. It looks like it'll come out to A=-3/7, but you should double-check that for me.

OpenStudy (anonymous):

what happened to the 2 in front of the e?

OpenStudy (anonymous):

i mean the 3

OpenStudy (anonymous):

That was from the RHS, this is all still solving for the LHS. The 3 is still there when you substitute and solve for A: \[(Ae^{5t})''-4(Ae^{5t})'-12(Ae^{5t})=3e^{5t}\]

OpenStudy (anonymous):

oh ok...so I guess for the sake of solving for A we leave it out?

OpenStudy (anonymous):

oh never mind. I see what you're doing

OpenStudy (anonymous):

Cool. Can you solve that for A and verify that it is -3/7?

OpenStudy (anonymous):

One more thing. So I it a general rule to say \[Y_P(t)=Ae^5t\] and we're using e^5t because that's in the original equation? ok now back to solving for A....

OpenStudy (anonymous):

Time to put it all together. The complementary solution is \[y_c=c_1e^{-2t}+c_2e^{6t}\] and the particular solution is \[Y_p=-\frac{3}{7}e^{5t}\] Altogether, the general solution is \[y(t)=c_1e^{-2t}+c_2e^{6t}-\frac{3}{7}e^{5t}\]

OpenStudy (anonymous):

and yes I got \[-\frac37\] for A

OpenStudy (anonymous):

It's a general rule to guess that y looks like Ae^{bt} because that is what was on the RHS. We know that you can differentiate Ae^{bt} as many times as you want and it'll still look pretty much like Ae^{bt}.

OpenStudy (anonymous):

got it. It's all coming together now

OpenStudy (anonymous):

If it was a different function, say, sin(x), then you'd want to guess something like sin(x) because the second derivative of sin(x) is -sin(x), so it only changed by a constant.

OpenStudy (anonymous):

Alright, so, now that we have the general solution, the only thing left is to apply the initial conditions to find c_1 and c_2.

OpenStudy (anonymous):

The initial conditions are in terms of y and y', so you'll have to differentiate our general y(t) to get that.

OpenStudy (anonymous):

ok \[\frac{18}7=c_1+c_2-\frac37\] and \[y'=-2c_1e^{-2t}+6c_2e^{6t}-\frac{15}7e^{5t}\]

OpenStudy (anonymous):

still working on it :)

OpenStudy (anonymous):

\[y'(0)=-\frac17=-2c_1+6c_2-\frac{15}{7}\]

OpenStudy (anonymous):

so now we have two equations and 2 unknowns...ok lets see

OpenStudy (anonymous):

\[c_2=41/14\]

OpenStudy (anonymous):

and c_1=1/7

OpenStudy (anonymous):

Hmm, that doesn't look right; let me double-check..

OpenStudy (anonymous):

Your equations look right, but they simplify to \[c_1+c_2=3\] \[-2c_1+6c_2=2\] Solving that system, c1=2 and c2=1

OpenStudy (anonymous):

what? How did you do that?

OpenStudy (anonymous):

c_1 + c_2 - 3/7 = 18/7 (from y_0) -2c_1 + 6c_2 - 15/7 = -1/7 (from y'_0) Those are the equations you got, right? Try solving that system again.

OpenStudy (anonymous):

I get -1 for c_2

OpenStudy (anonymous):

oh sorry I fault

OpenStudy (anonymous):

c_2 should be 2

OpenStudy (anonymous):

well I'll look at it again I guess =P gtg for now

OpenStudy (anonymous):



OpenStudy (anonymous):

Ok, so go through it all again and verify that the solution is \[y(t)=2e^{-2t}+e^{6t}-\frac{3}{7}e^{5t}\] Try taking the first and second derivatives of that and putting them back in the original equation to see if it's true and valid.